OpenStudy (anonymous):

Help pleasee.! help me get through it step by step pleasee 2|b+4|=48

OpenStudy (anonymous):

divide both sides by 2 :)

OpenStudy (anonymous):

|B+4| = 24

OpenStudy (anonymous):

then we get two equations due to modulo b+4 = 24 & b+4 = -24

OpenStudy (anonymous):

b = 24-4 & b = -24-4

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

can u solve it???

OpenStudy (anonymous):

welcm :)

OpenStudy (anonymous):

yea i can solve it . i just didnt know how to start it off

OpenStudy (anonymous):

:) n.p